OpenStudy (anonymous):

I'm having trouble finding the intersecting points of y=-3/4x+7(passes through the point (0,1)) and y=4/3x+1. Could someone help me with this?

OpenStudy (imstuck):

The intersection of these two lines is also called the solution of the system of equations. You could easily find the solution by making x/y charts for both equations, and whatever the identical point is in both charts is where they intersect.

OpenStudy (imstuck):

these lines are perpendicular, btw. Did you know that?

OpenStudy (imstuck):

The line y = -3/4x + 7 does not pass through the point (0,1).

OpenStudy (imstuck):

You could graph these lines in the same plane to see where they intersect, as well. Your choice. I'll do both for you so you can see, ok?

OpenStudy (imstuck):

Actually, the best way to find this solution is to solve them simultaneously. Use either the substitution method or the addition method.

OpenStudy (imstuck):

\[-\frac{ 3 }{4 }x+7=\frac{ 4 }{3 }x+1\]Solve that for x.

OpenStudy (imstuck):

When x = 2.88, y = 4.84 That is where they intersect.

OpenStudy (anonymous):

Thank you for responding! It doesn't go through (0,1)? That's what is says on my homework question. If you need it, I can do a screen cap of it. I tried x = 2.88 and y = 4.84 and it said it was wrong. Yes, I knew the lines were perpendicular.

OpenStudy (anonymous):

I guess I'll close this, talking to my professor about this. He may have worded it weird or put in the answers wrong.